# CSDO: Enhancing Efficiency and Success in Large-Scale Multi-Vehicle Trajectory Planning

### v1

@@ III-A2 Low-level planner
We directly adopt the spatiotemporal Hybrid A* from [2] as our low-level planner. Given a workspace $\mathcal{W}$ and static obstacle occupancy workspace $\mathcal{O}$, the higher priority agents’ trajectories $highTrajs$, a predefined start state $s_{i}$ and goal state $g_{i}$, the low-level path planner spatiotemporal hybrid A* (STHA*) will search a feasible trajectory. Compared with Hybrid A*, the STHA* add a time dimension to deal with the dynamic obstacles.
In the algorithm 2, it initiates by inserting the start node into open heap [Line 2]. Each node in open heap stores its state $z$ and timestamp t. The heap pops the node with minimum $f$ value, where $f=h+g$, $h$ represents the admissible heuristic function, similar to that used in Hybrid A*, and $g$ denotes the action cost from the start node to the current node. Generally, the greater the distance traveled and the more frequent reversing or changing direction, the higher the $g$ value. A node is popped from the heap for expansion [Lines 4-5]; if the expanded path is valid and collision-free, a solution can be obtained [Lines 6-9]. Otherwise, we iterate through the action set $Actions={FL,FS,FR,BL,BS,BR,Wait}$ to generate subsequent neighbor states of $n$ [Lines 10-11]. Invalid nodes are discarded [Lines 12-13], while valid ones calculated their $f$, $g$, and $h$ values [Lines 14-15]. If $n^{\prime}$ is not in the open heap, we push it into the heap [Lines 16-17]. Alternatively, if a node $n^{\prime}$ matches an existing heap node by timestamp, and state but offers a lower $g$ cost value, it replaces the existing one [Lines 18-21]. The whole process terminates with an empty solution if the heap empties[Line 22].
Input : $\mathcal{O},highTrajs,s_{i},g_{i}$
1 Open $\leftarrow$ new min heap sorted by node’s f value;
2 Push $(s_{i},t=0)$ into Open;
3 while $Open\neq\emptyset$ do
4       $Open\leftarrow$Open.pop();
5       $path_{toGoal}\leftarrow AnalyticExpand(n,g_{i})$ ;
6       if  $path_{toGoal}$ has no Collision with $highTrajs$ and $\mathcal{O}$ then
7             $path_{toN}\leftarrow$ Get Path By Finding $n$’s preceding nodes ;
8             $path\leftarrow path_{toN}$ + $path_{toGoal}$ ;
9             return $path$ ;
10
11      foreach  $action\in Actions$ do
12             $n^{\prime}\leftarrow$ Generate Next State from $n$ by $action$ ;
13             if  $n^{\prime}$ has collision with $highTrajs$ or $\mathcal{O}$ then
14                   continue ;
15                    // drop this node
16
17            $n^{\prime}.cost\leftarrow n.cost+cost(action)$ ;
18             $n^{\prime}.h\leftarrow n^{\prime}.cost+heuristic(n^{\prime},g_{i})$ ;
19             if  $n^{\prime}\notin$ Open then
20                   Push $n^{\prime}$ into Open ;
21
22            else
23                   $n_{same}^{\prime}\leftarrow$ same Discrete State as $n^{\prime}$ in Open;
24                   if $n^{\prime}.cost<n_{same}^{\prime}.cost$ then
25                         Replace the $n_{same}^{\prime}$ to $n^{\prime}$ in Open;
26
27
28
return $\emptyset$ ;
Algorithm 2 Spatiotemporal Hybrid A*
III-B Decentralized QP
As shown in Fig. 1, the decentralized QP takes an initial guess as input, then constructs safe corridors along them, and finally outputs trajectories for all agents.
@@ III-B Decentralized QP
III-B1 Initial Guess Interpolation
The initial guess, after a simple interpolation, may encounter three types of minor collisions, as illustrated in Fig. 4: type A: collisions between two vehicles, type B: collisions with obstacles, and type C: off-map states. For clarity, we denote $\bar{x}$ as the variable related to the interpolated initial guess. For instance, $\bar{z}^{(i)}_{t}$ is the state of agent $i$ in the timestamp $t$ in the dense initial guess.
To facilitate collision detection, we employ two uniformly distributed circles to cover the rectangular shape of the vehicle [13]. As illustrated in the Fig. 5, the circle centers are positioned at the quadrant points. The formulas for calculating the centers and radii of the two circles are as follows.
Figure 5: Generate separating plans for neighbor pairs.
$x^{F(i)}_{t}=x^{(i)}_{t}+L_{f2x}\cos{\theta^{(i)}_{t}},y^{F(i)}_{t}=y^{(i)}_{t}+L_{f2x}\sin{\theta^{(i)}_{t}},$ (8a)
$x^{R(i)}_{t}=x^{(i)}_{t}+L_{r2x}\cos{\theta^{(i)}_{t}},y^{R(i)}_{t}=y^{(i)}_{t}+L_{r2x}\sin{\theta^{(i)}_{t}},$ (8b)
@@ III-B1 Initial Guess Interpolation
where the $L_{f2x}=(3L_{F}-L_{B})/4$ is distance from the front disc center to the rear-axis center, $L_{r2x}=(L_{F}-3L_{B})/4$ is distance from the rear disc center to the rear-axis center, $L_{F}$ is the distance from rear axis to the front bumper, $L_{B}$ is distance from rear axis to the rear bumper. $Y^{F(i)}_{t}=[x^{F(i)}_{t},y^{F(i)}_{t}]^{T}$ is the center of the front circle. $Y^{R(i)}_{t}=[x^{R(i)}_{t},y^{R(i)}_{t}]^{T}$ refers to the rear circle’s center. $r_{v}$ is the radius of the circle. We denote $Y=[x^{F},y^{F},x^{R},y^{R}]^{T}$ as the vector of the circles positions. For any $i\in[M],0\leq t\leq\tau_{f}$, its linear form is as follows:
$Y_{t}^{(i)}=D_{t}^{(i)}\zeta_{t}^{(i)}+e_{t}^{(i)},$ (9a)
$D_{t}^{(i)}=\begin{bmatrix}1&0&-L_{f2x}\sin\bar{\theta}_{t}^{(i)}&0\\ 0&1&L_{f2x}\cos\bar{\theta}_{t}^{(i)}&0\\ 1&0&-L_{r2x}\sin\bar{\theta}_{t}^{(i)}&0\\ 0&1&L_{r2x}\cos\bar{\theta}_{t}^{(i)}&0\\ \end{bmatrix},\$ (9b)
$e_{t}^{(i)}=\begin{bmatrix}L_{f2x}(\cos{\bar{\theta}_{t}^{(i)}}+\bar{\theta}_{t}^{(i)}\sin\bar{\theta}_{t}^{(i)})\\ L_{f2x}(\sin{\bar{\theta}_{t}^{(i)}}-\bar{\theta}_{t}^{(i)}\cos\bar{\theta}_{t}^{(i)})\\ L_{r2x}(\cos{\bar{\theta}_{t}^{(i)}}+\bar{\theta}_{t}^{(i)}\sin\bar{\theta}_{t}^{(i)})\\ L_{r2x}(\sin{\bar{\theta}_{t}^{(i)}}-\bar{\theta}_{t}^{(i)}\cos\bar{\theta}_{t}^{(i)})\end{bmatrix}.\$ (9c)
III-B2 Neighbor Pair Searching
Given a distance threshold $R_{trust}$, we iterate through the plan to search pairs of agents $a^{(i)}$ and $a^{(j)}$ with a distance less than $2\sqrt{2}R_{trust}$ at the same timestamp. The distance function between two states $(\zeta^{(i)}_{t},\zeta^{(j)}_{t})$ is defined as follows:
$dist$ $(\zeta^{(i)}_{t},\zeta^{(j)}_{t})=\min(\left\|Y^{F(i)}_{t}-Y^{F(j)}_{t}\right\|,\left\|Y^{F(i)}_{t}-Y^{R(j)}_{t}\right\|,$ (10)
@@ III-B2 Neighbor Pair Searching
where the $Y^{F(i)}_{t}$, $Y^{R(i)}_{t}$, $Y^{F(j)}_{t}$, $Y^{R(j)}_{t}$ are the agent $a^{(i)}$’s front disc center, back disc center, agent $a^{(j)}$’s front disc center and rear disc center respectively. The search result, denoted as $NPairs$, comprises neighbor pairs and their corresponding timestamps. $NPairs=\{(a^{(i)},a^{(j)},t)|dist(\zeta^{(i)}_{t},\zeta^{(j)}_{t})<2\sqrt{2}R_{trust},\forall i,j\in[M],\forall 0\leq t\leq\tau_{f}\}$.
III-B3 Neighbor Pair Separation
After obtaining the neighbor pairs $NPairs$, we utilize the plane derived by the perpendicular bisector of the disc’s center as the constraint for mutual avoidance between vehicles. Each neighbor pair $(a^{(i)},a^{(j)},t)$ generates 4 separation planes {$C_{ffj}^{(i)},C_{frj}^{(i)},C_{rfj}^{(i)},C_{rrj}^{(i)}$} for agent $a^{(i)}$. Fig. 5 illustrates the process for agent $a^{(i)}$ and agent $a^{(j)}$—calculating perpendicular bisectors, offset by a distance of $r_{v}$ on both sides to obtain the separation half-plane $C_{ffj}^{(i)}$. Similarly, we generate the separation half-plane $C_{frj}^{(i)}$ for front-to-rear disc avoidance between agents $a^{(i)}$ and $a^{(j)}$, and obtain the separation half-planes $C_{rfj}^{(i)}$ and $C_{rrj}^{(i)}$ between the rear disc of agent $i$ and the discs of agent $a^{(j)}$.
$A_{c}^{(i)}Y^{(i)}\leq\overrightarrow{0},\forall i\in[M]$ (11)
where the $A_{c}^{(i)}$ denotes the corresponding half planes which are relevant to agent $a^{(i)}$ according to $NPairs$. When the initial guess is given, $A_{c}^{(i)}$ is a constant matrix.
@@ III-B3 Neighbor Pair Separation
$\left|Y^{(i)}_{t}-\bar{Y}^{(i)}_{t}\right|\leq R_{trust}\overrightarrow{1},\forall i\in[M],0\leq t\leq\tau_{f}.$ (12)
@@
For neighbor pairs, they are separated by the planes as in constraint (11). For non-neighbor pairs, they are separated by the variation range constraint (12).
@@
$A_{c}^{(i)}$ and $\bar{Y}^{(i)}_{t}$ are constants determined by the initial guess. For any two different vehicles $a^{(i)}$ and $a^{(j)}$, their variables $\zeta^{(i)}$ and $\zeta^{(j)}$ will not appear in the same inequality.
The above process decouples agents for distributed problem-solving. Without loss of generality, we describe the processing procedure for agent $a^{(i)}$ in the following steps.
@@ III-B4 Robust Corridor Construction
To handle the static obstacle avoidance constraint (5), we adapt the method from [23] to generate a corridor along its initial guess.
For clarity, ensuring that disc with radii $r_{v}$ does not go out of the map is equivalent to maintaining a distance of $r_{v}$ from the border. As in Fig. 6, we erode the map by a distance of $r_{v}$ to define the safety space enclosed by the dotted line. Similarly, we dilate the obstacles by a distance of $r_{v}$.
Note that our initial point may be in a colliding or out-of-map state, as previously mentioned in Fig. 4, causing the algorithm to immediately return an empty box. Therefore, we must relocate the initial point to a safe position before generating the box. In cases of a colliding state, we move the point outside of the grey circle. If it remains unsafe due to collisions with other obstacles, we gradually rotate it around the obstacle center until it becomes safe. For initial points that are out-of-map, we project them onto the map boundary. The corridor constraints can be summarized as follows,
$Y^{(i)}_{t,min}\leq A_{static,t}^{(i)}Y^{(i)_{t}}\leq Y^{(i)}_{t,max},\forall 0\leq t\leq\tau_{f},$ (13)
where $A_{static,t}^{(i)}$ denotes the generated corridor as in Fig. 6.
@@ III-B5 QP Formulation
Finally, we linearize the kinematic constraints. The objective function is set to minimize changes in velocity and steering wheel angle, aiming to smooth the trajectory. The kinematic constraints are linearized as follows:
$\zeta_{t+1}=A_{t}\zeta_{t}+B_{t}u_{t}+c_{t},$ (14a)
$A_{t}=\begin{bmatrix}1&0&-\bar{v}_{t}\sin\bar{\theta}_{t}*\Delta t&0\\ 0&1&-\bar{v}_{t}\cos\bar{\theta}_{t}*\Delta t&0\\ 0&0&1&\frac{\bar{v}_{t}*\Delta t}{L\cos^{2}\bar{\phi}}\\ 0&0&0&1\end{bmatrix},$ (14b)
$B_{t}=\begin{bmatrix}\cos\bar{\theta}_{t}*\Delta t&0\\ \sin\bar{\theta}_{t}*\Delta t&0\\ \frac{\tan\bar{\phi}*\Delta t}{L}&0\\ 0&\Delta t\end{bmatrix},$ (14c)
$c_{t}=[\bar{\theta}_{t}\bar{v}_{t}\sin\hat{\theta_{t}}\Delta t,-\bar{\theta}_{t}\bar{v}_{t}\cos\hat{\theta_{t}}\Delta t,-\frac{\bar{\phi}\Delta t}{L\cos^{2}{\phi_{t}}}]^{T},$ (14d)
where $(A_{t},B_{t},c_{t})$ are the associated coefficients at time $t$.
The objective function is as follows,
$J=\Sigma_{t}(\alpha_{v}(\Delta v_{t}^{(i)})^{2}+\alpha_{\omega}(\omega_{t}^{(i)})^{2}),$ (15)
where $\alpha_{v}$, $\alpha_{\omega}$ are the weighting parameters.
@@
$\min_{\zeta^{(i)},u^{(i)}}$ $J$ (16)
s.t. $\zeta_{0}^{(i)}=[s^{(i)T},0]^{T},\zeta_{\tau_{f}}^{(i)}=[g^{(i)T},0]^{T},$
$\left|v^{(i)}_{t}\right|\leq v_{max},\left|\omega^{(i)}_{t}\right|\leq\omega_{max},\forall 0\leq t<\tau_{f},$
$\left|\phi^{(i)}_{t}\right|\leq\phi_{max},\forall 0\leq t\leq\tau_{f},$
[…]

### v2

@@ III-A2 Low-level planner
We directly adopt the complete and optimal spatiotemporal Hybrid A* (STHA*) from [3] as our low-level planner. Compared with Hybrid A*, STHA* adds a time dimension to deal with the dynamic obstacles. Given a workspace $\mathcal{W}$ and static obstacle occupancy workspace $\mathcal{O}$, the higher priority agents’ trajectories $highTrajs$, a predefined start state $s_{i}$ and goal state $g_{i}$, STHA* will search the fastest trajectory when the solution space is not empty.
III-A3 Completeness and Optimality Analysis
p Refer to caption
Figure 4: PBS failed reason and well-formed scenarios.
Prioritized planning can fail to find a solution due to inappropriate priority orders [24]. The only issues are target blocking and run-over, as illustrated in Fig. 4. Target blocking occurs when a high-priority agent reaches its goal early and blocks a low-priority agent between its current position and goal. Run-over happens when a low-priority agent has no possible trajectory to avoid a high-priority agent. By utilizing tree search, PBS can explore all possible priority orders, making it P-complete. This capability allows PBS to greatly mitigate the above issues. Furthermore, in well-formed problems, all priority orders can lead to a feasible solution. So PBS can quickly find a feasible and even near-optimal solution. The key feature of well-formed problems is that agents can wait at their start and goal positions indefinitely without blocking other agents [20]. In practice, well-formed problems are very common, such as in intersection coordination[23]. Additionally, by changing the search strategy from depth-first search to best-first search, PBS can find a P-optimal solution, meaning it can find the best solution quality among the priorities, being optimal or near-optimal in practice [20].
III-B Decentralized SQP
As shown in Fig. 1, after inputting the initial guess, the separation planes are constructed to serve as inter-vehicle constraints. Then, multiple distributed SQP processes are employed to generate all the trajectories.
Notation. For clarity, the bar symbol represents the constant value. The subscript 0 denotes the constant associated with initial guess. For instance, $\bar{\theta}^{(i)}_{t,0}$ is the yaw angle of agent $a^{(i)}$ at timestamp $t$ in the interpolated initial guess. When there is no ambiguity, we omit the corresponding superscript $(i)$ or subscript $t$ when referring to all agents or any timestamp.
III-B1 Initial Guess Interpolation
Input initial guess $\bar{X}_{raw0}$. Output interpolated initial guess $\bar{X}_{0}$.
@@ III-B1 Initial Guess Interpolation
We interpolate each curve segment by inserting $n_{interp}$ points into each segment. So the time interval $\Delta t$ of adjacent points can be calculated as $\Delta t=\Delta S/((n_{interp}+1)v_{max})$. Afterwards, the initial guess may encounter three types of minor collisions, as illustrated in Fig. 5: type A: collisions between two vehicles, type B: collisions with obstacles, and type C: off-map states.
III-B2 Neighbor Pair Searching
Input Interpolated initial guess $\bar{X}_{0}$. Output Neighbor pairs $NPairs$.
To facilitate distance measuring for neighbor pair search and collision avoidance, we employ two uniformly distributed circles to cover the rectangular shape of the vehicle [15]. As illustrated in Fig. 6, the circle centers are positioned at the quadrant points. The formulas for calculating the centers and radii of the two circles are as follows.
Figure 6: Generate separating planes for neighbor pairs.
$x^{F}=x+L_{f2x}\cos{\theta},y^{F}=y+L_{f2x}\sin{\theta},$ (8a)
$x^{R}=x+L_{r2x}\cos{\theta},y^{R}=y+L_{r2x}\sin{\theta},$ (8b)
@@ III-B2 Neighbor Pair Searching
where the $L_{f2x}=(3L_{F}-L_{B})/4$ and $L_{r2x}=(L_{F}-3L_{B})/4$ are distance from the rear-axis center to the front and rear disc center, $L_{F}$ and $L_{R}$ are the distance from rear axis to the front bumper and rear bumper. $Y^{F}=[x^{F},y^{F}]^{T}$ and $Y^{R}=[x^{R},y^{R}]^{T}$ are the center of the front and rear circle. $r_{v}$ is the radius of the circle. We denote $Y=[x^{F},y^{F},x^{R},y^{R}]^{T}$ as the vector of the circles positions.
Given a distance threshold $R_{trust}$, we iterate through the plan to search pairs of agents $a^{(i)}$ and $a^{(j)}$ with a distance less than $2\sqrt{2}R_{trust}$ at the same timestamp. The distance function between two states $(z^{(i)}_{t},z^{(j)}_{t})$ is defined as follows:
$dist$ $(z^{(i)}_{t},z^{(j)}_{t})=\min(\left\|Y^{F(i)}_{t}-Y^{F(j)}_{t}\right\|,\left\|Y^{F(i)}_{t}-Y^{R(j)}_{t}\right\|,$ (9)
@@
where the $Y^{F(i)}_{t}$, $Y^{R(i)}_{t}$, $Y^{F(j)}_{t}$, $Y^{R(j)}_{t}$ are the agent $a^{(i)}$’s front disc center, back disc center, agent $a^{(j)}$’s front disc center and rear disc center respectively. The search result, denoted as $NPairs$, comprises neighbor pairs and their corresponding timestamps. $NPairs=\{(a^{(i)},a^{(j)},t)|dist(z^{(i)}_{t},z^{(j)}_{t})\leq 2\sqrt{2}R_{trust},\forall i,j\in[M],i\neq j,\forall 0\leq t\leq\tau_{f}\}$.
III-B3 Neighbor Pair Division
Input neighbor pairs $NPairs$. Output inter-collision avoidance constraints (10 - 11).
After obtaining the neighbor pairs $NPairs$, we utilize the plane derived by the perpendicular bisector of the disc’s center as the constraint for mutual avoidance between vehicles. Each neighbor pair $(a^{(i)},a^{(j)},t)$ generates 4 separation planes {$C_{ffj}^{(i)},C_{frj}^{(i)},C_{rfj}^{(i)},C_{rrj}^{(i)}$} for agent $a^{(i)}$. Fig. 6 illustrates the process. Calculating perpendicular bisectors, offset by a distance of $r_{v}$ to obtain the separation half-plane $C_{ffj}^{(i)}$. Similarly, we generate the front-to-rear, rear-to-front and rear-to-rear separation half-planes $C_{frj}^{(i)}$, $C_{rfj}^{(i)}$ and $C_{rrj}^{(i)}$, respectively.
$\bar{A}_{c,0}Y\leq\overrightarrow{0},$ (10)
where the $\bar{A}_{c,0}$ denotes the corresponding half planes. Notably, this method seamlessly adapts RSFC [25] under the assumption of two-circle approximation and discrete-time collision detection.
@@ III-B3 Neighbor Pair Division
$\left|Y-\bar{Y}_{0}\right|\leq R_{trust}\overrightarrow{1}.$ (11)
@@
Proof.
For neighbor pairs, they are separated by the planes as in constraint (10). For non-neighbor pairs, they are separated by the variation range constraint (11). ∎
@@
Proof.
$\bar{A}_{c,0}$ and $\bar{Y}_{0}$ are constants determined by the initial guess. For any two different vehicles $a^{(i)}$ and $a^{(j)}$, their variables $z^{(i)}$ and $z^{(j)}$ will not appear in the same inequality. ∎
The above process decouples agents for distributed problem-solving. Without loss of generality, we describe the processing procedure for agent $a^{(i)}$ in the following steps. This process is executed repeatedly until the stop criteria are met. Notation. The subscript $k$ represents the k-th iteration, with $k=0$ indicating the initial guess.
@@ III-B4 Robust Corridor Construction
Input obstacles $\mathcal{O}$ and $\bar{X}^{(i)}_{k}$. Output static collision avoidance constraints (12).
To handle the static obstacle avoidance constraint (5), we adapt the method from [26] to generate a corridor along the last iteration solution $\bar{X}_{k}^{(i)}$ of agent $a^{(i)}$.
For clarity, ensuring that a disc with radii $r_{v}$ does not go out of the map is equivalent to maintaining a distance of $r_{v}$ from the border. As in Fig. 7, we erode the map by a distance of $r_{v}$ to define the safety space enclosed by the dotted line. Similarly, we dilate the obstacles by a distance of $r_{v}$.
Note that our initial point may be in an out-of-map or colliding state, as previously mentioned in Fig. 5, causing the algorithm to immediately return an empty box. Therefore, we must relocate the initial point to a safe position before generating the box. For initial points that are out-of-map, we project them onto the map boundary. If the original or projected state collides with obstacles, we move the point outside of the grey circle. If it remains unsafe due to collisions with other obstacles, we gradually rotate it around the obstacle center until it becomes safe. The corridor constraints can be summarized as follows,
$\bar{Y}^{(i)}_{min,t,k}\leq\bar{A}_{static,t,k}^{(i)}Y^{(i)}_{t,k+1}\leq\bar{Y}^{(i)}_{max,t,k},\forall 0\leq t\leq\tau_{f},$ (12)
where $\bar{A}_{static,t,k}^{(i)}$ denotes the generated corridor as in Fig. 7.
@@ III-B5 QP Formulation
Input constraints (10-12) and last solution $\bar{X}^{(i)}_{k}$. Output Refine agent solution $\bar{X}^{(i)}_{k+1}$.
Finally, we linearize the kinematic constraints. This enables the smooth speed profile generation. The linearization error will be alleviated by the following sequential refinements. The objective function is set to minimize changes in velocity and steering wheel angle, aiming to smooth the trajectory. The objective function is as follows,
$J=\Sigma_{t}(\alpha_{v}(\Delta v_{t,k+1}^{(i)})^{2}+\alpha_{\omega}(\omega_{t,k+1}^{(i)})^{2}),$ (13)
where $\alpha_{v}$ and $\alpha_{\omega}$ are the weighting parameters.
The kinematic constraints are linearized as follows:
$z_{t+1,k+1}^{(i)}=\bar{A}_{t,k}^{(i)}z_{t,k+1}^{(i)}+\bar{B}_{t,k}^{(i)}u_{t,k+1}^{(i)}+\bar{c}_{t,k}^{(i)},\forall 0\leq t<\tau_{f}$ (14a)
$\bar{A}_{t,k}^{(i)}=\begin{bmatrix}1&0&-\bar{v}_{t,k}^{(i)}\sin\bar{\theta}_{t,k}^{(i)}*\Delta t&0\\ 0&1&-\bar{v}_{t,k}^{(i)}\cos\bar{\theta}_{t,k}^{(i)}*\Delta t&0\\ 0&0&1&\frac{\bar{v}_{t,k}^{(i)}*\Delta t}{L\cos^{2}\bar{\phi}_{t,k}^{(i)}}\\ 0&0&0&1\end{bmatrix},$ (14b)
$\bar{B}_{t,k}^{(i)}=\begin{bmatrix}\cos\bar{\theta}_{t,k}^{(i)}\Delta t&\sin\bar{\theta}_{t,k}^{(i)}\Delta t&\frac{\tan\bar{\phi}_{t,k}^{(i)}\Delta t}{L}&0\\ 0&0&0&\Delta t\end{bmatrix}^{T},$ (14c)
$\bar{c}_{t,k}^{(i)}=[\bar{\theta}_{t,k}^{(i)}\bar{v}_{t,k}^{(i)}\sin\hat{\theta_{t,k}^{(i)}}\Delta t,-\bar{\theta}_{t,k}^{(i)}\bar{v}_{t,k}^{(i)}\cos\hat{\theta_{t,k}^{(i)}}\Delta t,-\frac{\bar{\phi}_{t,k}^{(i)}\Delta t}{L\cos^{2}{\phi_{t,k}^{(i)}}}]^{T},$ (14d)
where $(\bar{A}_{t,k}^{(i)},\bar{B}_{t,k}^{(i)},\bar{c}_{t,k}^{(i)})$ are the associated coefficients.
To handle the non-linear calculation from state $z$ disc center positions $Y$, we need to linearize the Eq. (8), i.e.,
$Y_{t,k+1}^{(i)}=\bar{D}_{t,k}^{(i)}z_{t,k+1}^{(i)}+\bar{e}_{t,k}^{(i)},\forall 0\leq t\leq\tau_{f}$ (15a)
$\bar{D}_{t,k}^{(i)}=\begin{bmatrix}1&0&-L_{f2x}\sin\bar{\theta}_{t,k}^{(i)}&0\\ 0&1&L_{f2x}\cos\bar{\theta}_{t,k}^{(i)}&0\\ 1&0&-L_{r2x}\sin\bar{\theta}_{t,k}^{(i)}&0\\ 0&1&L_{r2x}\cos\bar{\theta}_{t,k}^{(i)}&0\\ \end{bmatrix},\$ (15b)
$\bar{e}_{t,k}^{(i)}=\begin{bmatrix}L_{f2x}(\cos{\bar{\theta}_{t,k}^{(i)}}+\bar{\theta}_{t,k}^{(i)}\sin\bar{\theta}_{t,k}^{(i)})\\ L_{f2x}(\sin{\bar{\theta}_{t,k}^{(i)}}-\bar{\theta}_{t,k}^{(i)}\cos\bar{\theta}_{t,k}^{(i)})\\ L_{r2x}(\cos{\bar{\theta}_{t,k}^{(i)}}+\bar{\theta}_{t,k}^{(i)}\sin\bar{\theta}_{t,k}^{(i)})\\ L_{r2x}(\sin{\bar{\theta}_{t,k}^{(i)}}-\bar{\theta}_{t,k}^{(i)}\cos\bar{\theta}_{t,k}^{(i)})\end{bmatrix}.\$ (15c)
@@
$\min_{{X}^{(i)}}$ $J$ (16)
s.t. $z_{0,k+1}^{(i)}=[s^{(i)T},0]^{T},z_{\tau_{f},k+1}^{(i)}=[g^{(i)T},0]^{T},$
$\left|v^{(i)}_{t,k+1}\right|\leq v_{max},\left|\omega^{(i)}_{t,k+1}\right|\leq\omega_{max},\forall 0\leq t<\tau_{f},$
$\left|\phi^{(i)}_{t,k+1}\right|\leq\phi_{max},\forall 0\leq t\leq\tau_{f},$
Constraints  (10)-(15)   related to $a^{(i)}$
III-B6 Stop Criteria
When the $||\bar{X}_{k+1}-\bar{X}_{k}||_{2}$ is less than the given threshold, or the plan is feasible, we stop the iteration.